# MAD-Sherlock: Multi-Agent Debates for Out-of-Context Misinformation Detection

### v1

@@ Abstract
One of the most challenging forms of misinformation involves the out-of-context (OOC) use of images paired with misleading text, creating false narratives. Existing AI-driven detection systems lack explainability and require expensive finetuning. We address these issues with MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection. MAD-Sherlock introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making. Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning. Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that MAD-Sherlock boosts performance for both experts and non-experts. These results position MAD-Sherlock as a powerful tool for autonomous and citizen intelligence applications.
@@ 1 Introduction
Our growing dependence on online channels for news and social networking has been complemented by a surge in exploits of digital misinformation (Aslett et al., 2024; Hasher et al., 1977; Brashier & Marsh, 2020). While many manipulation techniques pose serious threats, one of the most prevalent methods for creating fake online content is the out-of-context (OOC) use of images (pbs, ). This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise. Indeed, the problem of OOC misinformation detection requires a complex understanding of the relationship between the text and image and the ability to identify when they do not go together. Identifying these minute inconsistencies is a time-consuming and high-effort task for humans. A study by Sultan et al. (2022) shows that time pressure reduces the ability of human beings to detect misinformation effectively, further adding to the scalability issues in human expert detection.
Therefore, attention has turned to AI-driven tools that can help human experts recognise instances of OOC image-based misinformation at scale. Unfortunately, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), which target detecting manipulations such as PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated (or manipulated) fake images called Deepfakes (mit, ), rely on spotting artifacts from image or text tampering. In contrast, OOC detection demands cross-contextual reasoning, as the deception arises from the misalignment between the legitimate image and its falsely associated textual content.
@@
Figure 1: Overview of MAD-Sherlock: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not. After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
@@
In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3). Specifically, our framework MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allow for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b). In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of MAD-Sherlock as an interactive tool for human experts. To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
We perform a comprehensive empirical evaluation of our method (see Section 4), including the study of multiple debate configurations. To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance. We find that MAD-Sherlock outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies. We identify both access to external information retrieval and complete freedom of opinion as key ingredients to MAD-Sherlock’s performance. Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
@@ 2 Related Work
Recent work has focused on using joint image-text representations to classify an instance as OOC. Aneja et al. (2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar. They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption. During inference, they use the semantic similarity between the two captions to classify them as OOC or not. The increased reliance on textual content limits the capabilities of this approach. This work also does not provide explanations for model predictions and is, therefore not interpretable. Moreover, this method works for image caption pairs where captions have information about objects in the image. This is not always the case with news articles (our domain of application), where captions can often just be related to the main content of an article rather than precisely describing the objects in the image. Appendix  A.2 shows an example of the same.
@@
Abdelnabi et al. (2022) present the Consistency Checking Network (CCN) in which they emulate different aspects of human reasoning across modalities for misinformation detection. This method uses evidence related to the image-text pair aggregated from the Internet. The CCN consists of memory networks to assess the consistency of the image-caption pair against the retrieved evidence and a CLIP (Radford et al. (2021)) component to evaluate the consistency between the image and caption pair. The use of external evidence to better inform model decisions is an important idea and also explains the superior classification performance of CCN when compared to other methods. This method also lacks the explainability component.
@@
Zhang et al. (2024) extend the neural symbolic method (Yi et al. (2019); Zhu et al. (2022)) to propose an interpretable cross-modal misinformation detection model to provide supporting evidence for the output prediction. They use symbolic graphs based on the Abstract Meaning Representation (Banarescu et al. (2013)) of textual and visual information to detect OOC image use. Zhou et al. (2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks are used to learn features of text and visual news representations. Their representations and relationships are jointly learned and used to predict fake news. Wang et al. (2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated. EANN uses adversarial training to learn multi-modal features independent of news events. These methods require pretraining from scratch and, therefore, don’t benefit from the advanced reasoning capabilities and world knowledge of large pretrained models.
@@
Shalabi et al. (2023) use synthetic multi-modal data to establish the authenticity of image-text pairs. They use BLIP-2 (Li et al. (2023b)) to generate a caption for the original image and Stable Diffusion (Rombach et al. (2022)) to generate an image for the given original caption. This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text. This method relies on synthetic multi-modal data generation, which not only adds an additional computational overhead but also increases dependence on often unreliable synthetically generated data. Therefore, this method can suffer from issues related to generation models, including potential biases that these models may possess. This method also lacks interpretability.
@@
Sniffer (Qi et al. (2024)) is the closest to our work. It uses the InstructBLIP (Dai et al. (2023)) model to detect OOC image use and provide an explanation for its prediction. It makes use of internal and external knowledge using entity extraction APIs and image-based web searches. Information from all the sources is given to an LLM to predict and explain if an image has been used OOC. Sniffer only uses basic textual information, such as news article titles from websites, to form its external knowledge base. It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
Lin et al. (2024) use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful. These explanations are then used to train a small language model as a judge to determine whether the image and text that make up the meme are actually harmful. This work does not allow agents to have flexibility of opinion. There are always two agents, and each one is provided a stance to defend. Moreover, a judge decides the final outcome of the debate and needs to be trained on data from the debate. This method also does not benefit from external retrieval, and therefore, the debating agents are not aware of the crucial external context related to the input. Finally, this work is related to harmful meme detection and does not concern the problem of misinformation detection in the news domain, which likely requires more intricate contextual analysis, including of external context.
@@ 3 Methodology
We present an explainable misinformation detection system, MAD-Sherlock, which jointly predicts and explains instances of misinformation. Figure 1 illustrates our approach. To the best of our knowledge, all prior work except Qi et al. (2024) provide predictions without explanations, and no prior work uses multiple models to approach this problem. We present a novel methodology that involves multiple multi-modal models debating against each other in order to decide if an image-text pair is misinformation or not. In this work, we aim to answer the question:
Can debating multi-modal models, when equipped with external context, be used to solve the problem of explainable misinformation detection by picking up on minute contextual inconsistencies?
We use detailed external information to inform the model’s predictions through the external retrieval module, which utilises reverse image-based search in order to provide the agents with external real-world context related to the image-text pair. We carry out our experiments with the GPT-4o (OpenAI ) model to achieve state-of-the-art performance on the misinformation detection task while also providing detailed and coherent explanations for the predictions. We achieve this without any domain-specific fine-tuning, thus ensuring easier and faster generalization to other domains in addition to low computational overhead.
@@ 3.1 Debate Modelling
Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways. We explore multiple debating strategies to structure the conversation between agents. Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s). We observe this facilitates more involved and detailed discussions among the models.
@@
We define an asynchronous debating strategy in which models wait for the other participants’ responses before generating their own. Figure 2 (a) and (b) show synchronous and asynchronous debating structures, respectively. While synchronous debates, where all participants speak at once, can be faster and computationally more efficient, we opt for an asynchronous setting where each model response is based on previous responses of other models. We observe that this method of structuring the debate works better since models are able to pick up on contextual ambiguities in their responses in a more organised and structured way which is crucial to the process of misinformation detection. An important point to note for this setup is the way we structure model prompts. The debating models are not aware that they are debating other AI agents. The prompts are structured such that each debating model believes it is talking to a human.
@@
We also experiment with an asynchronous debate setup with a judge. Figure 2 (c) shows this setup. In this setup, models participate in an asynchronous debate as usual however, the final decision is made by a judge at the end of the debate. Models are incentivised to structure their arguments in a way that makes them most convincing to the judge. We structure this debate configuration similar to Khan et al. (2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
@@
Improving on the actor-skeptic method, in this setup, we allow all agents to act as actors and skeptics. Models are tasked with not only generating their own responses but also disambiguation queries to refine further or refute the other agents’ responses. These disambiguation queries are then used to search the Internet to obtain information to refine model outputs further. We are the first to propose this debate setup and while it does not achieve the best results in this work, we believe future research can greatly benefit by refining this setup further.
We empirically identify the best debate set up to be the asynchronous debate where the model believes it is debating against a human and not an AI agent.
@@ 3.2 Prompt Engineering
The debate structure is substantiated through prompt engineering. Figure 1 shows that the first stage of our method requires for each AI agent to generate an independent response to whether the given image-text pair is misinformation. Each agent must take into account the external context related to the image obtained through the external information retrieval module. Specifications of the various prompts used in this work can be found in Appendix A.3. An initial prompt provides the agent with a summary of the news articles related to the image and, based on it, asks the agent to classify the image-text pair as misinformation or not. The prompt asks the agent to focus on certain details in the image, such as watermarks, flags, etc. We observe that images used in news articles often contain minute yet crucial details which can be used to inform the final decision about whether the image actually belongs to the news articles. Therefore, prompting the agent to pay special attention to these details further helps detect inconsistencies.
Once the agents have formed independent opinions about the image-text pair, they must then participate in a debate. While the same prompt can be used for each debate round, we provide a different prompt for the first round to allow each agent to understand the changing nature of the conversation. Responses from other agents are provided as a part of the prompt, and the agent is asked to agree or disagree. The agent must also clearly state the reasoning process behind its argument. The prompt further requires the agent to identify ambiguities in the other agents’ reasoning. This allows the agent to closely analyse the responses from other agents and use them to inform its own decision.
A separate prompt is then used to facilitate all rounds of the debate following round one. It takes into account the other agent’s suggestions from the previous round and, after refining its own response, asks the agent to point out any inconsistencies in this new response. We note that in such a scenario, agents are prone to simply agreeing with each other and repeating the other agent’s response. This is especially the case when agents believe they are conversing with a human. However, this tendency of the agents to easily give up their own opinions and simply agree with the other participants does not facilitate an advantageous debate, and the agents do not discover any new information related to the image-text pair. Therefore, the debate prompt also contains clear and explicit instructions asking the agent not to simply agree with the presented response unless it has an acceptable reason to do so. We find this helps the agents develop stronger stances and reluctance to blind agreement.
@@ 3.3 External Information Retrieval
Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al. (2023)). This summary is then passed to the debating agents as a part of the initial prompt.
Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame). Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al. (2022)) to supplement external information related to an image-caption pair. However, we find this information lacking in detail since it is limited to the title of a news article. Agents can greatly benefit from the knowledge of the entire news article and its content rather than just the title when making a decision about whether a given image-caption pair, when considered in the context of the news article, is misinformation. To this end, we propose our own external information retrieval module. We observe a significant improvement in accuracy after incorporation of the external information retrieval module into the pipeline. The module is implemented in two stages:

### v2

# LLM-Consensus: Multi-Agent Debate for Visual Misinformation Detection

@@ Abstract
One of the most challenging forms of misinformation involves the out-of-context (OOC) use of images paired with misleading text, creating false narratives. Existing AI-driven detection systems lack explainability and require expensive finetuning. We address these issues with LLM-Consensus, a multi-agent debate system for OOC misinformation detection. LLM-Consensus introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making. Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning. Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that LLM-Consensus boosts performance for both experts and non-experts. These results position LLM-Consensus as a powerful tool for autonomous and citizen intelligence applications.
misinformation detection,out-of-context image use,LLMs,multimodal models,multi-agent debate,safety
@@ 1 Introduction
Our growing dependence on online channels for news and social networking has been complemented by a surge in exploits of digital misinformation (Aslett et al., 2024; Hasher et al., 1977; Brashier & Marsh, 2020). While many manipulation techniques pose serious threats, one of the most prevalent methods for creating fake online content is the out-of-context (OOC) use of images (pbs, ). This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise.
OOC misinformation detection requires a nuanced understanding of the relationship between text and images and the ability to identify when they are misaligned. Detecting these subtle inconsistencies is time-consuming for humans, and (Sultan et al., 2022) shows that time pressure further reduces detection accuracy, exacerbating scalability issues.
As a result, AI-driven tools have gained attention for scaling detection efforts. However, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), designed to detect manipulations like PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated Deepfakes (mit, ), focus on spotting artifacts from tampering. In contrast, OOC detection requires cross-contextual reasoning, as the deception stems from the misalignment of legitimate images with misleading text.
Figure 1: Overview of LLM-Consensus: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not. After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
@@
In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3). Specifically, our framework LLM-Consensus frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allows for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b). In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of LLM-Consensus as an interactive tool for human experts. To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
We perform a comprehensive empirical evaluation of our method (see Section 4), including the study of multiple debate configurations. To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance. We find that LLM-Consensus outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies. We identify both access to external information retrieval and complete freedom of opinion as key ingredients to the performance of LLM-Consensus. Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
@@ 2 Related Work
Recent work has focused on using joint image-text representations to classify an instance as OOC. (Aneja et al., 2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar. They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption. During inference, they use the semantic similarity between the two captions to classify them as OOC or not. The increased reliance on textual content limits the capabilities of this approach. This work also does not provide explanations for model predictions and is, therefore not interpretable. Moreover, this method works for image caption pairs where captions have information about objects in the image. This is not always the case with news articles (our domain of application), where captions can often just be related to the main content of an article rather than precisely describing the objects in the image. Appendix  A.2 shows an example of the same.
@@
(Abdelnabi et al., 2022) introduce the Consistency Checking Network (CCN), which models aspects of human reasoning across modalities for misinformation detection using external evidence aggregated from the Internet. The CCN employs memory networks to evaluate the consistency of image-caption pairs against retrieved evidence and a CLIP component (Radford et al., 2021) to assess the alignment between the image and caption. The use of external evidence enhances CCN’s classification performance compared to other methods. However, it lacks an explainability component, providing only binary decisions without further explanation.
@@
(Zhang et al., 2024) extend the neural symbolic method (Yi et al., 2019; Zhu et al., 2022) to develop an interpretable cross-modal misinformation detection model that provides supporting evidence for its predictions. They use symbolic graphs based on Abstract Meaning Representation (Banarescu et al., 2013) to detect out-of-context image use. Similarly, (Zhou et al., 2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks jointly learn and analyze features and relationships between text and visual news representations to predict fake news. (Wang et al., 2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated. EANN uses adversarial training to learn multi-modal features independent of news events. These methods require pretraining from scratch and, therefore, don’t benefit from the advanced reasoning capabilities and world knowledge of large pretrained models.
@@
Shalabi et al. (2023) use synthetic multi-modal data to establish the authenticity of image-text pairs. They use BLIP-2 (Li et al., 2023b) to generate a caption for the original image and Stable Diffusion (Rombach et al., 2022) to generate an image for the given original caption. This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text. This method relies on synthetic multi-modal data generation, which not only adds an additional computational overhead but also increases dependence on often unreliable synthetically generated data. Therefore, this method can suffer from issues related to generation models, including potential biases that these models may possess. This method also lacks interpretability.
@@
Sniffer (Qi et al., 2024) is the closest to our work. It uses the InstructBLIP (Dai et al., 2023) model to detect OOC image use and provide an explanation for its prediction. It makes use of internal and external knowledge using entity extraction APIs and image-based web searches. Information from all the sources is given to an LLM to predict and explain if an image has been used OOC. Sniffer only uses basic textual information, such as news article titles from websites, to form its external knowledge base. It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
@@ 3 Methodology
We present LLM-Consensus, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1). Unlike prior work, which largely provides predictions without explanations, our approach uses multiple multi-modal models debating to determine whether an image-text pair constitutes misinformation. We address the question:
Can debating multi-modal models, equipped with external context, detect misinformation by identifying subtle contextual inconsistencies?
Our external retrieval module uses reverse image search to provide agents with real-world context, enhancing their predictions. Using the GPT-4o (OpenAI, ) model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning. This ensures faster generalization to new domains with minimal computational overhead.
@@ 3.1 Debate Modelling
Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways. We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments. Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s). We observe this facilitates more involved and detailed discussions among the models.
@@
One of the core setups we test in our experiments is an asynchronous debating strategy, where models wait for the responses of other participants before generating their own. Figure 2 (a) and (b) illustrate the differences between synchronous and asynchronous debate structures. While synchronous debates, where all participants respond simultaneously, can be faster and more computationally efficient, we find the asynchronous setup more effective. This structure allows models to better identify contextual ambiguities in a more organized and systematic manner, which is crucial for misinformation detection. A key aspect of this setup is the design of model prompts: the debating models are not aware that they are interacting with other AI agents but are instead led to believe they are debating with a human.
@@
We also experiment with an asynchronous debate setup with a judge. Figure 2 (c) shows this setup. In this setup, models participate in an asynchronous debate as usual however, the final decision is made by a judge at the end of the debate. Models are incentivised to structure their arguments in a way that makes them most convincing to the judge. We structure this debate configuration similar to (Khan et al., 2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
@@
Building on the actor-skeptic method, we allow all agents to act as both actors and skeptics. Models generate their own responses and disambiguation queries to refine or challenge other agents’ outputs. These queries are used to retrieve additional information from the Internet, further improving model responses.
@@ 3.2 Prompt Engineering
The debate structure is enabled through prompt engineering. As shown in Figure 1, the first stage involves each AI agent independently determining whether the image-text pair is misinformation, using context from the external retrieval module (details in Appendix A.3). The initial prompt summarizes related news articles and instructs the agent to focus on key image details, like watermarks and flags, to identify inconsistencies.
After forming initial opinions, the agents engage in a debate. The first round uses a unique prompt to help agents adapt to the discussion, while subsequent rounds use a standard prompt incorporating suggestions from previous rounds. Agents must agree or disagree with others’ responses while identifying ambiguities and refining their reasoning.
To prevent blind agreement, the prompt explicitly instructs agents to provide valid reasoning before accepting others’ responses. This promotes stronger, independent stances and enhances the discovery of new information.
@@ 3.3 External Information Retrieval
Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis, ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al., 2023). This summary is then passed to the debating agents as a part of the initial prompt.
A model’s world knowledge is limited by its training data and time frame, but external retrieval allows access to up-to-date information beyond these constraints. Previous work relies on external datasets (Abdelnabi et al., 2022), but these often provide only news article titles, which lack sufficient detail. Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame). Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al., 2022) to supplement external information related to an image-caption pair. However, we find this information lacking in detail since it is limited to the title of a news article. Full access to the article content can significantly enhance an agent’s ability to assess whether an image-caption pair constitutes misinformation. To address this, we propose a dedicated external information retrieval module, which improves accuracy when integrated into the pipeline. The module operates in two stages: